The perimeter of a rectangle is 34 inches, and its area is 70 square inches. Find the length and width of the rectangle.
Answer by checkley79(3341)

P=2L+2W
P=34
34=2L+2W
2L=-2W+34
L=-2W/2+34/2
L=-W+17
A=LW
70=(-w+17)W
70=-W^2+17W
W^2-17W+70=0
(W-10)(W-7)=0
W-10=0
W=10 ANS. & LENGTH IS 7.
W-7=0
W=7 ANS. & LENGTH IS 10.
PROOF:
34=2*7+2*7
34=14+20
34=34
70=7*10
70=70